A loaded dogsled has mass of 400 kg and is being pulled across a horizonal, packed snow surface at a velocity of 4.0 m/s (N). Suddenly the harness separates from the sled. If the coefficient of kinetic friction for the sled on the snow is 0.0500, how far will the sled coast before stopping?


1
Expert's answer
2021-10-27T08:48:40-0400

We can find the distance that the sled will coast before stopping from the work-kinetic energy theorem:


"\Delta KE=W_{fr},""\dfrac{1}{2}mv^2-\dfrac{1}{2}mv_0^2=-F_{fr}d,""-\dfrac{1}{2}mv_0^2=-\mu_kmgd,""d=\dfrac{v_0^2}{2\mu_kg}=\dfrac{(4.0\ \dfrac{m}{s})^2}{2\times0.05\times9.8\ \dfrac{m}{s^2}}=16\ m."